OpenStudy (anonymous):

x^2+kx-19 find alll the values of k so that the trinomial can be factored using intergers

OpenStudy (anonymous):

lets see what the answer is

myininaya (myininaya):

-19 -19 and 1 are factors of -19 so we need k=-19+1=-18

myininaya (myininaya):

or -19 19 and -1 are factors of -19 so k needs to be 19-1=18

myininaya (myininaya):

two options

OpenStudy (anonymous):

lol

myininaya (myininaya):

x^2-18x-19 (x-19)(x+1) ------------- x^2+18x-19 (x+19)(x-1)

OpenStudy (anonymous):

k could alo be -18

OpenStudy (anonymous):

also !!

myininaya (myininaya):

yes it cool only be -18 or 18

OpenStudy (anonymous):

Are you sure k=18 works?

myininaya (myininaya):

(x+19)(x-1)

OpenStudy (anonymous):

yes (x + 19)(x - 1) = x^2 + 18x - 19